# Magneto: Combining Small and Large Language Models for Schema Matching

### v1

@@ 2.2. Magneto: Overview
Magneto consists of two main components, a candidate retriever and a reranker, which use small and large language models to efficiently derive accurate matches. The intuition is to first apply a cheaper approach to retrieve and filter candidates so that a more sophisticated (and expensive) methods can be used to accurately identify the correct matches from a smaller candidate set. By doing so, we reduce the overall cost of schema matching and, at the same time, attain high accuracy. Figure 2 shows one variant of Magneto that uses a fine-tuned language model as the retriever, and a large language model as the reranker. The use of language models enables Magneto to serve as general solution that supports schema matching across multiple domains.
Candidate Retriever. The candidate retriever leverages SLMs to generate a ranked list of potential matches from the target table for each input column. It uses column embeddings to estimate column pair similarities (Cappuzzo et al., 2020; Fan et al., 2023). SLMs are a good choice for this step given their ability to capture semantic similarity and efficiency.
@@
Pre-trained (generic) SLMs such as BERT (Devlin, 2018), RoBERTa (Liu, 2019), and MPNet (Song et al., 2020), can struggle with nuanced syntactic differences and lack contextual knowledge for complex schema matching tasks involving domains not represented in their training data. To perform complex tasks effectively, these models need to be fine tuned. We propose to leverage LLMs to automatically generate the data needed to fine tune SLMs. We show that LLMs can help derive high-quality training data that reflect instances of variability for semantically similar columns that arise in real data. This approach, which we describe in Section 3, improves the robustness of the SLM-based retriever, making it possible for it handle complex matches without requiring human-labeled training data. We refer to Magneto configurations that use this approach using the label ft. Note that Magneto-ft only invokes LLMs once, during the training phase for a given domain. Once trained, the fine-tuned SLM can perform multiple inferences efficiently without making calls to LLMs.
@@
Match Reranker. While fine-tuning improves the performance of SLMs, it is not sufficient to handle schema matching tasks that include domains and heterogeneity not covered during the fine tuning. To improve generalizability, Magneto uses LLMs as rerankers to refine the candidates identified by the SLM-based retriever. This approach, which we refer to as Magneto-llm, enhances accuracy by leveraging carefully designed prompts and techniques that enable the LLM to judiciously assess matches and discern subtle semantic nuances, which are challenging for the SLM to detect independently. As discussed in Section 4, Magneto-llm also strives to minimize costs by optimizing LLM resources.
Varying Retrievers and Rerankers. The architecture of Magneto allows the combination of different alternatives for retrievers and rerankers. This is useful for studying the tradeoffs of various approaches and to derive solutions that match the accuracy requirements and computational constraints for different applications.
@@
For the candidate retriever, we can combine traditional matching techniques with the embedding strategies. For example, when columns in a match have the same name, module minor variations in case and punctuation, we can assign a perfect similarity score of 1.0 to the match. During our experiments, this simple technique consistently improved overall accuracy, albeit slightly, in most cases.
@@
We implemented an alternative match reranker, the bipartite-graph reranker, which adapts the a filtering technique from (Melnik et al., 2002). This approach offers an algorithmic alternative for reranking, making it particularly suitable for scenarios where LLMs are unavailable or where strict runtime constraints must be met. This solution combines all ranked match candidates of each source column into one global ranked list and transforms it into an undirected bipartite graph where the nodes represent columns. There are two (disjoint) node sets consisting of the columns in the source and target tables (respectively). A match is represented as an edge in the graph, with a corresponding score as its edge weight. To identify the best match for the set of attributes in the source column, it uses the algorithm from (Crouse, 2016), which solves the assignment problem in polynomial time and is efficient even for large graphs. We prioritize the matches in the assignment by placing the corresponding edges at the top of the list. To account for the remaining matches, we move them to the bottom of the list while preserving their relative order..
@@ 3. Using LLMs to Fine-Tune SLMs
SLMs for Candidate Match Retrieval. Pre-trained small language models (SLMs) have been used to help with schema-matching related tasks by encoding semantic information from column names and values into dense vector representations–embeddings (Tu et al., 2023; Zhang et al., 2023; Fan et al., 2023; Cong et al., 2023; Dong et al., 2023; Du et al., 2024). To identify matches, embeddings of source and target columns can be compared using cosine similarity: high similarity scores indicate a higher likelihood that the columns match.
SLMs are pre-trained on and work well for general natural language tasks, but their ability to interpret tabular data is limited. These models must often be fine-tuned to handle tasks that involve tables. However, fine-tuning approaches require large amounts of labeled data for training (Tu et al., 2023; Dong et al., 2023). Unfortunately, this is impractical in many scenarios such as the integration of biomedical data (Example 1.1), for which training data is hard to obtain.
In the absence of training data, it is possible to apply augmentation techniques to automatically generate variations of data to be used as positive examples. For example, given a column, different versions can be generated by shuffling rows, sampling values, and applying perturbations to values (Fan et al., 2023; Dong et al., 2023). However, these variations may not fully capture the heterogeneity found in real data.
We introduce a new method that leverages LLMs to generate training data and present a pipeline to fine tune SLMs for schema matching (Section 3.2). Another important consideration in the use of SLMs, that we discuss this in Section 3.1, is how to represent tables and their columns.
Figure 3. LLM-Powered Fine-tuning Pipeline
3.1. Column Serialization and Value Sampling
@@
Given column $A$ with column type $type(A)$ and sample values $v_{1},v_{2},\dots,v_{m}$, we consider the following approaches:
@@
where the [CLS] token is used as a special symbol to indicate the start of a column, and [SEP] separates the components within a serialized column.
$\mathcal{S}_{\texttt{default}}$ is a variant of a widely-adopted serialization strategy (Suhara et al., 2022). Instead of employing commas to separate sample values, we utilize the [SEP] token. This modification prevents the model from interpreting the values as an ordered text sequence, thereby treating them as unordered, discrete features. This change is crucial in schema matching as it emphasizes the structured and independent nature of the data in each column, enabling more accurate alignment of similar but non-identical schemata.
$\mathcal{S}_{\texttt{verbose}}$ is an extension of $\mathcal{S}_{\texttt{default}}$ where prefixes are added to delineate each component (i.e., column name and values) and provide additional context for the SLM.
$\mathcal{S}_{\texttt{repeat}}$ repeats the column name multiple times to reinforce its importance, nudging the model to prioritize column names. It attempts to mitigate attention drift in SLMs, which can disproportionately focus on values rather than on column names, especially in zero-shot settings. Specifically, we chose $k=5$ because it strikes a balance between reinforcing column names without overwhelming the model with redundancy.
@@
Priority Sampling for Column Values. To efficiently retrieve candidate matches, we generate embeddings by including a sample of column values in the serialized column representation. Different sampling techniques have been explored for various table management tasks, such as column type annotation (Feuer et al., 2024) and table union search (Fan et al., 2023). A common approach is to use weighted sampling and assign higher weights to more frequent values. We adopt this approach and incorporate coordination into the sampling process with priority sampling (Daliri et al., 2024). Priority sampling was used for inner product sketching and aimed to maximize the likelihood of selecting corresponding values across vectors, emphasizing those with larger magnitudes.
In our setting, priority sampling is adapted to increase the probability of sampling similar values during the fine-tuning and inference stages. Specifically, we use a random seed $s$ to select a uniformly random hash function $h:\{1,\dots,N\}\to[0,1]$, where $N$ represents the maximum number of unique values across all columns. For each value $v_{i}$, we compute a rank:
$R_{i}=\frac{\text{freq}(v_{i})}{h(v_{i})},$
@@ 3.2. LLM-Powered Fine Tuning
By fine-tuning a model, our goal is for it to learn column representations so that its embeddings can spatially cluster related columns. This spatial arrangement helps the retrieval of semantically related representations using cosine similarity. With this goal in mind, the basic intuition behind our training methodology is to minimize the distance between embeddings of identical or semantically-related columns while maximizing the distance between distinct columns.
@@
Synthetic columns are derived from an input (anchor) column as follows: 1) The anchor name and a sample of its values to the LLM using the structured prompt depicted in Figure 4, and 2) the LLM outputs columns that are semantically similar to the anchor. Examples of synthetic columns are shown in Table 2. This method establishes a class of columns considered matches (derived from the same anchor); columns within the same class serve as positive examples and those from different classes as negative examples.
@@
We combine the LLM-based augmentation with other structure-based augmentation augmentation methods (struct-aug) such as random sampling and shuffling of values, and minor perturbations to the column name, including random character replacements or deletions (Fan et al., 2023). These techniques inject variability while preserving syntactic and structural similarity to the original column. By using both augmentation strategies, the model can learn to identify matches with different characteristics.
Note that we restrict this fine-tuning and training process to the target table columns to avoid inaccuracies. Applying this process to source tables with unknown true matches to target tables could mistakenly classify them as negative examples, introducing errors into the embedding space.
Triplet Loss and Online Triplet Mining. To leverage the generated synthetic columns, we implement a contrastive learning framework using triplet loss and online triplet mining (Schroff et al., 2015). Triplet loss improves upon previous contrastive loss functions by more directly optimizing for relative similarities, which, as we discovered in our experiments, yields more precise schema alignments.
In the training pipeline, each triplet consists of an anchor column $a$, a positive column $p$, and a negative column $n$; $a$ and $p$ are semantically related (i.e., they belong to the same class), while $n$ is drawn from a different class. The goal is to learn an embedding space such that the distance between the $a$ and $p$ is minimized while the distance between $a$ and $n$ is maximized. Specifically, we incorporate Batch Hard triplet loss (Schroff et al., 2015) to refine the learning of discriminative embeddings for column classification. Given $P$ classes and $K$ columns per class, the loss function is defined as follows:
@@
Here, $a$ is the anchor column embedding from a specific class $i$, $p$ is the positive column embedding within the same class as the anchor, and $n$ is the negative column embedding from a different class. The function $D$ denotes the distance metric, cosine similarity in our case, which measures the similarity between embeddings. The margin $m$ is a hyperparameter that ensures a minimum distance between the positive and negative embeddings relative to the anchor, thereby enhancing the separability between classes. This loss function iterates over all classes ($P$) and multiple examples per class ($K$) within each training batch, optimizing the embeddings to emphasize inter-class distinctions and intra-class similarities.
@@
    Hardest Positives: Triplets where the distance between the anchor and the positive is maximized: $D(a,p)$ is maximized such that $D(a,p)<D(a,n)$
@@
    Hardest Negatives: Triplets where the distance between the anchor and the negative is minimized, yet still greater than the anchor-positive distance plus the margin: $D(a,n)$ is minimized but still $D(a,n)>D(a,p)+m$
@@
    Semi-Hard Negatives: These are negatives that are not the closest to the anchor but are still closer than the positive plus the margin: $D(a,p)<D(a,n)<D(a,p)+m$
These conditions help the model not settle for easy examples and instead learn to distinguish subtle differences, developing more robust and discriminative features to distinguish columns. By combining triplet loss and online triplet mining, our model can effectively utilize the synthetic columns to enhance the discriminative power of the embeddings for schema matching.
Model Selection for Schema Matching. To select the optimal fine-tuned model during training, we must define an effective validation metric specifically tailored for effective schema matching tasks. Our selection process uses the measures described in Section 2, mean reciprocal rank (MRR) and recall at ground truth (Recall@GT), to evaluate model performance on synthetic data. Since no ground-truth data is available, we rely on synthetic datasets to simulate real-world scenarios. These metrics are specifically chosen to ensure that the model identifies correct matches and ranks them in a manner that reflects their true relevance.
The intuition behind using these metrics, rather than a traditional loss metric, lies in their ability to reflect the practical requirements of schema matching. Loss metrics typically provide a general sense of model error but do not differentiate between the types of errors that are more consequential for the task. In schema matching, failing to rank a highly relevant match at the top (as evaluated by MRR) or missing a relevant match entirely (as evaluated by recall) can be more detrimental than the errors captured by average loss.
@@ 4. LLMs as Rerankers
SLMs can serve as efficient retrievers, but they may fail to capture complex semantic relationships. Moreover, models fine-tuned with LLM-derived data are inherently domain-specific. To mitigate these limitations, Magneto incorporates a Large Language Model (LLM) as a reranker. Unlike existing methods that adjust match rankings based on heuristic rules and similarity metrics scores (Madhavan et al., 2001; Melnik et al., 2002; Do and Rahm, 2002; Gal, 2006; Radwan et al., 2009; Alserafi et al., 2020), Magneto leverages the knowledge accumulated by the LLM to complement the initial retrieval and improve the overall performance of SM.
A natural question arises: Why not just use an LLM for SM? While there are approaches that employ LLMs for SM (Li et al., 2024a; Parciak et al., 2024), their applicability in scenarios that involve many or large tables is limited due to high computational costs and challenges related to context windows (Challenge 2, Section 1). We posit that LLMs must be used judiciously and designed Magneto accordingly.
@@
Input Source Column:[Serialized Source Column]
Target Columns:[Serialized Candidate Columns]
@@
Scoring-Oriented Prompt Design. The Scoring-Oriented Instruction is a critical component of the prompt that aligns the model with the schema matching task. Specifically, this instruction involves sending one source column and the top $k$ candidate target columns retrieved by the SLM to the LLM. Previous works, such as (Li et al., 2024a) and (Parciak et al., 2024), typically employ table-wise prompts that ask LLMs to generate a ranked list without providing individual scores, which poses scalability challenges when entire tables are involved. Furthermore, column-wise ranking without scoring prevents effective table-wise comparisons, such as evaluating the recall of matches against the ground truth (recall@GT) for each pair of source and target columns.
To address these limitations, we propose a novel prompt design that requires the model to assign scores from 0.00 to 1.00 to each column pair, rather than merely producing a ranked list. This not only facilitates direct comparisons across different pairs but also allows the model to adopt a holistic view by considering all $k$ target columns simultaneously during scoring. Additionally, our serialization of the column data clarifies to the model that a text sequence comprising a column name followed by its values represents a column in a relational table. This assists in more accurately scoring matches based not only on the semantic and contextual relevance but also on the structural characteristics of the column data.
Few-Shot In-Context Learning. By providing a few examples of a task to an LLM, few-shot learning can lead to significant improvement in LLM performance (Brown, 2020; Dong et al., 2022). This method has been shown effective in various data management tasks, enabling robust performance without the need for extensive fine-tuning (Hegselmann et al., 2023; Shraga et al., 2020).
In the One-shot Example strategy, we provide a single example of the schema matching task, which outputs a ranked list with scores. This approach clarifies both the objective of the task and expected output format, as well as establishes a uniform scoring standard, enhancing the comparability of matches across all source and target column pairs. Utilizing only one example helps maintain a lightweight prompt structure, crucial for minimizing input and output token counts. The number of tokens directly impacts runtime and cost, and it can also influence accuracy (Li et al., 2024b).
Optimization of Model Cost. For the implementation of Magneto, we use GPT-4o-mini, a low-cost LLM, which has shown good performance for our prompts. To ensure reliability, we attempt to parse the model’s output up to three times. If parsing fails after these attempts, we revert to using the embedding scores as a fallback mechanism. During testing, output-related issues were infrequent; however, they occurred more commonly with larger tables and in scenarios lacking a candidate retriever.

### v2

@@ 2.2. Magneto: Overview
Magneto first applies a cheaper approach to retrieve and filter candidates so that a more sophisticated method can accurately identify the correct matches from a smaller candidate set. Figure 2 shows one variant of Magneto that uses a fine-tuned language model as the retriever and a large language model as the reranker.
@@
Candidate Retriever. The candidate retriever leverages an SLM to generate a ranked list of potential matches from the target table for each input column. It uses column embeddings to estimate column pair similarities (Cappuzzo et al., 2020; Fan et al., 2023). SLMs are a good choice for this step given their ability to capture semantic similarity and efficiency. General-purpose pre-trained SLMs such as BERT (Devlin, 2018), RoBERTa (Liu, 2019), and MPNet (Song et al., 2020), may struggle with syntactic differences and lack contextual knowledge for domains absent in their training data. To perform complex tasks effectively, these models need to be fine-tuned. We propose to leverage LLMs to automatically generate the data needed to fine tune SLMs. We show that LLMs can help derive high-quality training data that reflect instances of variability for semantically similar columns that arise in real data. This approach, described in Section 3, improves the robustness of the SLM-based retriever, making it possible for it to handle complex matches without requiring human-labeled training data. We refer to Magneto configurations that use this approach using the label ft. Note that Magneto-ft invokes LLMs only during the offline training phase for a given domain. Once trained, the fine-tuned SLM can perform multiple inferences efficiently without making calls to LLMs.
@@
Match Reranker. While fine-tuning improves SLM performance, it is not sufficient to handle schema matching tasks involving domains and heterogeneity unseen during the fine tuning. To improve generalizability, Magneto uses LLMs as rerankers to refine the candidates identified by the SLM-based retriever. This approach, which we refer to as Magneto-llm, enhances accuracy by leveraging carefully designed prompts and techniques that enable the LLM to judiciously assess matches and discern subtle semantic nuances, which are challenging for the SLM to detect independently. As discussed in Section 4, Magneto-llm also reduces LLM costs.
@@
Varying Retrievers and Rerankers. The architecture of Magneto allows the combination of different alternatives for retrievers and rerankers. We can combine traditional matching techniques with the embedding strategies for the candidate retriever. For example, when columns in a match have the same name, modulo minor variations in case and punctuation, we assign a perfect similarity score of 1.0. During our experiments, this simple technique consistently improved overall accuracy, albeit slightly, in most cases.
@@
We also implemented bipartite-graph reranker as an algorithmic alternative that adapts the filtering technique from Melnik et al. (Melnik et al., 2002). This approach, which we refer as Magneto-bp, is particularly suitable for scenarios where LLMs are unavailable during inference or where strict runtime constraints must be met. The algorithm combines all ranked match candidates across source columns into a single global list and transforms it into an undirected bipartite graph. There are two (disjoint) node sets consisting of the columns in the source and target tables (respectively). The graph contains two disjoint node sets representing the source and target table columns, respectively, with potential matches represented as weighted edges between nodes. Edge weights correspond to match confidence scores. To identify the best match for the set of attributes in the source column, it uses the algorithm from (Crouse, 2016), which solves the assignment problem in polynomial time and scales to large graphs. The final ranking prioritizes matches selected by the assignment algorithm at the top of the list, while unselected matches are placed at the bottom in their original relative order.
@@ 3. Using LLMs to Fine-Tune SLMs
Pre-trained small language models (SLMs) have been used to help with schema-matching related tasks by encoding semantic information from column names and values into dense vector representations–embeddings (Tu et al., 2023; Zhang et al., 2023; Fan et al., 2023; Cong et al., 2023; Dong et al., 2023; Du et al., 2024). To identify matches, embeddings of source and target columns can be compared using cosine similarity: high similarity scores indicate a higher likelihood that the columns match. SLMs work well for general natural language tasks, but their ability to interpret tabular data is limited. These models must often be fine-tuned to handle tasks that involve tables. However, fine-tuning approaches require large amounts of labeled data for training (Tu et al., 2023; Dong et al., 2023). Unfortunately, this is impractical in many scenarios such as the integration of biomedical data (Example 1.1), for which training data is hard to obtain.
In the absence of training data, it is possible to apply augmentation techniques to automatically generate variations of data to be used as positive examples. For example, given a column, different versions can be generated by shuffling rows, sampling values, and applying perturbations to values (Fan et al., 2023; Dong et al., 2023). However, these variations may not fully capture the heterogeneity found in real data. We introduce a new method that leverages LLMs to generate training data and present a pipeline to fine-tune SLMs for schema matching (Section 3.2). Another key consideration is how to represent tables and their columns, which we discuss in Section 3.1.
3.1. Value Sampling and Column Serialization
Priority Sampling for Column Values. We generate embeddings to retrieve candidate matches by including a sample of column values in the column representation. Sampling strategies have been adopted for data management tasks, including column type annotation (Feuer et al., 2024) and table union search (Fan et al., 2023). A common approach is to use weighted sampling and assign higher weights to more frequent values. We adopt this approach and incorporate coordination into the sampling process with priority sampling (Daliri et al., 2024). For inner product sketching, priority sampling maximizes the likelihood of selecting corresponding values across vectors by emphasizing elements with larger magnitudes.
In our setting, priority sampling is adapted to optimize the selection of column values. This approach not only prioritizes frequent values, which are statistically more representative of the column domain, but it also increases the likelihood of identifying shared values across different columns that act as inter-column anchors. These anchors enhance similarity detection in SLMs, which are sensitive to token co-occurrence patterns learned during pre-training. Specifically, we use a random seed $s$ to select a uniformly random hash function $h:\{1,\dots,N\}\to[0,1]$, where $N$ represents the maximum number of unique values across all columns. For each value $v_{i}$, we compute a rank:
$R_{i}=\frac{\text{freq}(v_{i})}{h(v_{i})},$
@@ 3.1. Value Sampling and Column Serialization
Given column $A$ with column type $type(A)$ and sample values $v_{1},v_{2},\dots,v_{m}$, we consider the following serialization approaches:
@@
where the [CLS] token is a special symbol that indicates the start of a column, and [SEP] separates column components.
$\mathcal{S}_{\texttt{default}}$ is a variant of a widely-adopted serialization strategy (Suhara et al., 2022; Yin et al., 2020). Instead of employing commas to separate sample values, we utilize the [SEP] token. This modification prevents the model from interpreting the values as an ordered text sequence, thereby treating them as unordered, discrete features. This change is crucial in schema matching as it emphasizes the structured and independent nature of the data in each column.
$\mathcal{S}_{\texttt{verbose}}$ is an extension of $\mathcal{S}_{\texttt{default}}$ where prefixes are added to delineate each component and provide additional context for the SLM. By explicitly tagging each data segment, it helps models better contextualize the information, leading to improved interpretability and alignment accuracy in schema matching tasks.
$\mathcal{S}_{\texttt{repeat}}$ repeats the column name multiple times to reinforce its importance, nudging the model to prioritize column names. This strategy is inspired by findings in attention-based neural network research where repetition can enhance item salience (Vaswani et al., 2017). It attempts to mitigate attention drift in SLMs, which can disproportionately focus on later tokens or values rather than column names, especially in zero-shot settings where the model is not optimized to learn the importance of different column components. For our experiments, we set $k=5$ to strike a balance between reinforcing column names without overwhelming the model with redundancy.
@@ 3.2. LLM-Powered Fine Tuning
By fine-tuning, we aim to learn column representations whose embeddings cluster spatially to reflect semantic relationships. This spatial arrangement enables efficient retrieval of related columns via cosine similarity. Our training methodology thus minimizes the distance between embeddings of identical or semantically related columns while maximizing separation between distance ones.
@@
Synthetic columns are derived from an input (anchor) column as follows: 1) The anchor name and a sample of its values are given to the LLM using the structured prompt depicted in Figure 4, and 2) the LLM outputs columns that are semantically similar to the anchor. Examples of synthetic columns are shown in Table 2. This method establishes a class of columns considered matches (derived from the same anchor); columns within the same class serve as positive examples and those from different classes as negative examples.
@@
We combine the LLM-based augmentation with other structure-based augmentation methods (struct-aug) such as random sampling and shuffling of values, and minor perturbations to the column name, including random character replacements or deletions (Fan et al., 2023). These techniques inject variability while preserving syntactic and structural similarity to the original column. By using both augmentation strategies, the model can learn to identify matches with different characteristics. We restrict this fine-tuning and training process to the target table columns to avoid inaccuracies. Applying this process to source tables with unknown true matches to target tables could mistakenly classify them as negative examples, introducing errors into the embedding space.
Triplet Loss and Online Triplet Mining. To leverage the generated synthetic columns, we implement a contrastive learning framework using triplet loss and online triplet mining (Schroff et al., 2015). Unlike standard contrastive losses, triplet loss directly optimizes relative similarities, improving schema alignment precision.
Each training triplet includes an anchor column $a$, a same-class positive $p$, and a different-class negative $n$. The model learns to minimize $D(a,p)$ and maximize $D(a,n)$, where $D$ is cosine distance. Specifically, we incorporate Batch Hard triplet loss (Schroff et al., 2015) which is defined as follows:
@@
Here, $m$ is a margin hyperparameter that ensures a minimum distance between the positive and negative embeddings relative to the anchor, thereby enhancing the separability between classes. The loss iterates over $P$ classes and $K$ columns within each training batch, optimizing the embeddings to emphasize inter-class distinctions and intra-class similarities.
@@
    Hard Negatives: Closest negative to the anchor that violates the margin constraint: $n^{*}=\arg\min_{n}D(a,n)\text{ where }D(a,n)<D(a,p)+m$
@@
    Semi-Hard Negatives: Negatives farther than the positive but within the margin: $D(a,p)<D(a,n)<D(a,p)+m$
@@
    Hard Positives: Farthest positive from the anchor within its class: $p^{*}=\arg\max_{p}D(a,p)$
These conditions help the model not settle for easy examples and instead learn to distinguish subtle differences, developing more robust and discriminative features to distinguish columns. Together, triplet loss and mining improve embedding discriminability for schema matching.
Model Selection for Schema Matching. To select the optimal fine-tuned model during training, we must define an effective validation metric specifically tailored for schema matching tasks. Our selection process uses the measures described in Section 2.1 to evaluate model performance on synthetic data. Since no ground-truth data is available, we rely on synthetic datasets to simulate real-world scenarios. These metrics are specifically chosen to ensure that the model identifies correct matches and ranks them in a manner that reflects their true relevance.
@@ 4. LLMs as Rerankers
SLMs can serve as efficient retrievers, but they may fail to capture complex semantic relationships. Moreover, models fine-tuned with LLM-derived data are inherently domain-specific. To mitigate these limitations, Magneto incorporates a Large Language Model (LLM) as a reranker. Unlike existing methods that adjust match rankings based on heuristic and similarity metrics (Madhavan et al., 2001; Melnik et al., 2002; Do and Rahm, 2002; Gal, 2006; Radwan et al., 2009; Alserafi et al., 2020), Magneto leverages LLM understanding to complement initial retrieval and improve overall schema matching performance.
A natural question arises: Why not just use an LLM for schema matching? While there are approaches that employ LLMs for schema matching (Li et al., 2024a; Parciak et al., 2024), their applicability in scenarios that involve many or large tables is limited due to high computational costs and challenges related to context windows (Challenge 2, Section 1). We also empirically demonstrate that LLM-only approaches underperform an SLM-LLM combination for large target schemas (Figure 11). We posit that LLMs must be used judiciously and designed Magneto accordingly.
@@
Input Candidate Column::[Serialized Source Column]
Target Schemas:[Serialized Target Columns]
@@
Scoring-Oriented Prompt Design. The Scoring-Oriented Instruction aligns the model with schema matching by sending one source column and the top $k$ target candidates from the SLM to the LLM. Prior works (Li et al., 2024a; Parciak et al., 2024) use table-wise prompts that rank columns without individual scores, limiting scalability. Column-wise ranking without scoring also hinders table-wise comparisons, such as recall@GT evaluation. To address these limitations, we propose a novel prompt design that requires the model to assign scores from 0.00 to 1.00 to each column pair, rather than merely producing a ranked list. This facilitates direct comparisons across different pairs and allows the model to adopt a holistic view by considering all $k$ target columns simultaneously during scoring. Additionally, our serialization of the column data clarifies to the model that a text sequence comprising a column name followed by its values represents a column in a relational table. This allows more accurately scoring matches based not only on the semantic and contextual relevance but also on the structural characteristics of the data.
Few-Shot In-Context Learning. By providing a few examples of a task to an LLM, few-shot learning can lead to significant improvement in LLM performance (Brown et al., 2020; Dong et al., 2022). This method has been shown effective in various data management tasks, enabling robust performance without the need for extensive fine-tuning (Hegselmann et al., 2023; Shraga et al., 2020).
In the One-shot Example strategy, we provide a single example of the schema matching task, which outputs a ranked list with scores. This approach clarifies both the objective of the task and expected output format, and establishes a uniform scoring standard, enhancing the comparability of matches across all source and target column pairs. Utilizing only one example helps maintain a lightweight prompt structure, crucial for minimizing input and output token counts. The number of tokens directly impacts runtime and cost, and it can also influence accuracy (Li et al., 2024b).
Optimization of Model Cost. To ensure reliability, we attempt to parse the LLM’s output up to three times. If parsing fails after these attempts, we revert to using the embedding scores as a fallback mechanism. During testing, output-related issues were infrequent; however, they occurred more commonly with larger tables and in scenarios lacking a candidate retriever.